*Production volume for product A was estimated at 1000 units.

*Only 80% production was achieved.

*Each unit of product A requires 0.5 hours at an hourly rate of N$8.50.

*80% production volume was produced at 45 minutes per unit at an hourly rate of N$8.00

*Buys raw materials from local suppliers at N$3.50 per kilograms.

*Two kilograms were used per unit.

*Total monthly budgeted manufacturing overheads costs amounted to N$950

*Actual monthly manufacturing overheads cost was N$1200

*Lungameni Enterprises production manager view that manufacturing overheads costs be absorbed on basis of direct labour cost. There was no opening inventory and all units produced were sold.


Required:

1.For the information above calculate the gross profit / loss.

2.Any recommendation for Lungameni Enterprise production manager




1
Expert's answer
2022-04-11T16:53:11-0400

1) Gross profit/loss can be found as PQ-TC, where P is the price for product A, Q is the production volume and TC is total cost. The information about price of product A is absent above, so we cannot find the gross profit/loss. We can find only total cost: 1000×0.8×0.8×8×0.75+1000×0.8×0.2×0.5×8.5+1000×0.8×2×3.5+1200=3840+680+5600+1200=11320

2) The enterprise should strive for the estimated parametres of production, because it provides lower total cost